If in a triangle, the sum of two angles is equal to the third angle then the triangle is called – what? Right/acute/obtuse?
and How?

Answer by richard1234(7193)

If the sum of two of the angles is equal to the third angle, then the third angle must be 90 degrees. This is because if we let the angles be A,B,C with A+B = C, we have A+B+C = 180, replace A+B with C to get 2C = 180, C = 90. Hence the triangle is a right triangle.